2. Find the gradient of the curve y =4 sin 2x + 3cos 1/2x, at x = TT/3


1
Expert's answer
2021-05-07T10:10:30-0400

Solution:

"y=4\sin 2x+3\cos (1/2\ x)"

On differentiating both sides w.r.t "x",

"\frac{dy}{dx}=4\cos 2x.(2)+3(-\sin (\frac12 \ x)).(\frac12)
\\=8\cos 2x-\frac32 (\sin \frac12x)"

Put "x=\frac{\pi}3"

"\frac{dy}{dx}=8\cos 2(\frac{\pi}3)-\frac32 (\sin \frac12.\frac{\pi}3)
\\=8\cos (\frac{2\pi}3)-\frac32 (\sin \frac{\pi}6)
\\=8(\frac{-1}{2})-\frac{3}{2}(\frac{1}{2})
\\=-4.75"

So, the gradient of "y" at "x=\frac{\pi}3" is "-4.75".